# LocMoE: A Low-Overhead MoE for Large Language Model Training

### v2

@@ 1 Introduction
  • Orthogonal gating weight with Global Average Pooling (GAP) layer. The GAP layer is adopted in gating values computation. It provides a natural way to perform class activation mapping and reduce computational costs. Above all, the orthogonality of gating weight facilitates the explicit decisions of the router.
@@ 3.2 MoE With Local Routing Strategy
Figure 4: Difference between feature extraction via the dense layer and the GAP layer.
where $i^{*}$ is the index of the most appropriate expert, and $\mathcal{G}_{m,E_{i}}=\mathrm{ReLU}(\omega_{i}\cdot x_{m}+\epsilon_{i})$. The token would be sent to the Top-1 expert with the largest expert score screened by Softmax. To reduce the parameter scale and the computational overhead, the gating value is obtained via the GAP layer instead of the dense layer [?]. The feature extraction with the GAP layer is delineated in Figure 4. It can be regarded as the dense layer with the fixed weight $\omega_{i}$:
@@
where $d$ denotes the dimension of activation. Notably, the gating weights of the GAP layer are orthogonal. From a perspective of semantic, irrelevant tokens are inclined to be routed to experts of different domains, which is conducive to convergence and accuracy [?]. Besides, the GAP layer has greater computation efficiency.
@@
Suppose $i_{j}$ denotes the expert to which the token $j$ is routed. On account of the spherical symmetry, the probabilities for $i_{j}=i^{{}^{\prime}}$ are equivalent for all experts under the conditions of orthogonal gating weights (brought from the GAP layer). That is, $P\{i_{j}=i^{{}^{\prime}}\}=\frac{1}{n}$.
@@ 5 Conclusion
In this paper, we propose a low overhead structure named LocMoE to relieve the performance bottleneck of existing MoE. The modifications mainly revolve around the mechanism of token assignment. The locality loss, which can be delineated as the distribution difference of token assignments, is proposed to promote locality computation on the premise of load balance. We also provide the theoretical demonstration for the lower bound of the expert capacity to achieve the same effect by training fewer tokens. To meet the assumption of orthogonal gating weight, the GAP layer is adopted instead of the dense layer to calculate the gating values, and it can also reduce the overhead of computation. Incorporating group-wise All-to-All and communication overlapping features, the elapsed time of communication is further reduced. The experiments are performed on Ascend clusters with 64, 128, and 256 910A NPUs. Compared with current state-of-the-art MoEs, the performance improvement of training is up to 22.24%. Evaluating multiple NLP tasks, it is detected that the interactive capability of our model is also enhanced. From the results that explore the relationship between the scale of expert capacity and the token features, we find that the dataset construction still needs to be improved. In future work, we will further organize the multilingual corpora from more fields.

### v3

@@ 1 Introduction
  • Orthogonal gating weight with Grouped Average Pooling (GrAP) layer. The GrAP layer is adopted in gating values computation. It provides a natural way to perform class activation mapping and reduce computational costs. Above all, the orthogonality of gating weight facilitates the explicit decisions of the router.
@@ 3.2 MoE With Local Routing Strategy
Figure 4: Difference between feature extraction via the dense layer and the GrAP layer.
where $i^{*}$ is the index of the most appropriate expert, and $\mathcal{G}_{m,E_{i}}=\mathrm{ReLU}(\omega_{i}\cdot x_{m}+\epsilon_{i})$. The token would be sent to the Top-1 expert with the largest expert score screened by Softmax. To reduce the parameter scale and the computational overhead, the gating value is obtained via the GrAP layer instead of the dense layer [?]. The feature extraction with the GrAP layer is delineated in Figure 4. It can be regarded as the dense layer with the fixed weight $\omega_{i}$:
@@
where $d$ denotes the dimension of activation. Notably, the gating weights of the GrAP layer are orthogonal. From a perspective of semantic, irrelevant tokens are inclined to be routed to experts of different domains, which is conducive to convergence and accuracy [?]. Besides, the GrAP layer has greater computation efficiency.
@@
Suppose $i_{j}$ denotes the expert to which the token $j$ is routed. On account of the spherical symmetry, the probabilities for $i_{j}=i^{{}^{\prime}}$ are equivalent for all experts under the conditions of orthogonal gating weights (brought from the GrAP layer). That is, $P\{i_{j}=i^{{}^{\prime}}\}=\frac{1}{n}$.
@@ 5 Conclusion
In this paper, we propose a low overhead structure named LocMoE to relieve the performance bottleneck of existing MoE. The modifications mainly revolve around the mechanism of token assignment. The locality loss, which can be delineated as the distribution difference of token assignments, is proposed to promote locality computation on the premise of load balance. We also provide the theoretical demonstration for the lower bound of the expert capacity to achieve the same effect by training fewer tokens. To meet the assumption of orthogonal gating weight, the GrAP layer is adopted instead of the dense layer to calculate the gating values, and it can also reduce the overhead of computation. Incorporating group-wise All-to-All and communication overlapping features, the elapsed time of communication is further reduced. The experiments are performed on Ascend clusters with 64, 128, and 256 910A NPUs. Compared with current state-of-the-art MoEs, the performance improvement of training is up to 22.24%. Evaluating multiple NLP tasks, it is detected that the interactive capability of our model is also enhanced. From the results that explore the relationship between the scale of expert capacity and the token features, we find that the dataset construction still needs to be improved. In future work, we will further organize the multilingual corpora from more fields.